OpenStudy (anonymous):

The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger?

OpenStudy (anonymous):

42 56 182 194

OpenStudy (anonymous):

so basically you keep adding 18 and 8together until you reach the 2,520 amount

OpenStudy (anonymous):

okay???

OpenStudy (anonymous):

@SeaTurtle113 please help!!!!

OpenStudy (seaturtle113):

x+y=224 people x=18 and Under age y=+18 age 8x+12y=$2520 With this, and I believe you know how to do it with this info, you should solve with x.

OpenStudy (anonymous):

224 = x+y 2520 = 8x+12y x being the visitors under 18, y being the ones over 18.

OpenStudy (anonymous):

Oh I'm late, darn it. xD.

OpenStudy (anonymous):

It's basically a system of equations. :) You can substitute and stuff

OpenStudy (anonymous):

okay yea I got that but my problem is solving for the answer

OpenStudy (anonymous):

No wait, not substitution.. elimination :$ but it looks like sea did it for you.

OpenStudy (anonymous):

@SeaTurtle113 thats not the right answer I put the possibilites at the top. Sorry

OpenStudy (seaturtle113):

I know... that was an accident.

OpenStudy (anonymous):

oh sorry lol.

OpenStudy (seaturtle113):

Haha! No sorry! My bad!

OpenStudy (anonymous):

no problem but can you help me solve this one pleaseeeee?????

OpenStudy (seaturtle113):

I'm trying!! It's been awhile since I've done this.

OpenStudy (anonymous):

i'm sorry!!!!

OpenStudy (seaturtle113):

lol You're good.... :)))

OpenStudy (seaturtle113):

I'm sorry I can't figure it out, but I know someone who can... when do you need this answer?

OpenStudy (anonymous):

now!!!!!!

OpenStudy (seaturtle113):

Ugh.. sorry. :(

OpenStudy (anonymous):

no problem :p